Please help w/ the following:
Find all the real and complex zeros of the polynomial x^4+8x^3+8x^2-72x-153.
Thank you.

Answer by stanbon(75887)

Find all the real and complex zeros of the polynomial x^4+8x^3+8x^2-72x-153.
----------
Graphing I find that x=3 and x=-3 are zeroes.
Dividing the polynomial by x^2-9 I find the
quotient is x^2+8x+17.
Using the quadratic formula I get:
x=[-8+-sqrt(64-4(17)]/2
x=[-8+-sqrt(-4)]/2
x=[-4+-i]
Real zeros are x=3 and x=-3
Complex zeros are x=-4+i and x=-4-i
Cheers,
Stan H.